# LV-XAttn: Distributed Cross-Attention for Long Visual Inputs in Multimodal Large Language Models

### v1

@@ 4.1 Experimental Setup
Model Setup We evaluate our methods and baselines on 5 models shown in Table 2. Following their default configuration, each frame is encoded into 729 visual tokens for the mPLUG-Owl3 models and 64 visual tokens for the OpenFlamingo models. This implies that given the same amount of memory capacity, we can fit more frames to OpenFlamingo models than mPLUG-Owl3 models.
@@
Baselines For our method, we use LV-XAttn for the cross-attention layers and Ring Attention for the LM blocks. Our primary baseline is the setup where Ring Attention is used for both the cross-attention layers and LM blocks. We apply our activation recomputation technique to both of these settings for enabling longer context length. We also compare against Deepspeed-Ulysses (Jacobs et al., 2024), which employs sequence parallelism for non-attention layers and head parallelism for attention layers. All methods uses Flash Attention.
@@ 4.2 Comparison with Ring Attention
Table 3: Per iteration wall-clock time (in seconds) on 16 A100 80GB GPUs with Ring Attention and LV-XAttn. “CA” represents the time spent on cross-attention operations. As $S_{Q}$ doubles, the cross-attention speedup nearly halves because the runtime for Ring Attention, which is communication-bound, remains constant, while the runtime for LV-XAttention, which is computation-bound, doubles. On the other hand, as $S_{KV}$ doubles, both communication and computation also double, so the speedup remains roughly the same.
@@
length count CA (s) Total (s) CA (s) Total (s) CA Total
@@
Table 3 shows the per iteration time of 5 models using LV-XAttn and Ring Attention on 16 A100 80GB GPUs. For the mPLUG-Owl3 models, LV-XAttn speeds up the cross-attention operation by 7.04 – 15.32$\times$. Since the cross-attention operation accounts for the majority of the total iteration time when using Ring Attention, this reduction results in a significant total iteration speedup of 3.3 – 5.58$\times$. For the OpenFlamingo models, which process a larger number of frames and thus have longer text lengths (due to the inclusion of a special token <image> per frame) and larger $S_{Q}$, the speedup is less pronounced, LV-XAttn achieves 1.47 – 2.16$\times$ speedup on the cross-attention operation and 1.16 – 1.92$\times$ speedup on the total iteration time. Additionally, OpenFlamingo-3b, with denser cross-attention layers, spends a larger portion of its time in cross-attention compared to OpenFlamingo-9b when using Ring Attention. Consequently, the speedup in cross-attention translates to a more substantial end-to-end speedup for OpenFlamingo-3b.
Table 4 shows the same experiment on 8 A30 24GB GPUs. We have smaller text lengths and fewer frames due to the smaller memory capacity. In this setup, the speedup for cross-attention operation is greater than that on 16 A100 GPUs: 20.6 – 45.85$\times$ for the mPLUG-Owl3 models and 3.97 – 7.2$\times$ for the OpenFlamingo models. This is due to smaller query block sizes $\frac{S_{Q}}{n}$ (shorter computations favors computation-bound LV-XAttn) and slower interconnect bandwidth (longer communication hurts communication-bound Ring Attention), as shown in Table 1. However, the larger cross-attention speedups do not translate into a larger total speedup, as the portion of time spent on cross-attention layers decreases due to slower self-attention layers in LM blocks (caused by the slower interconnect). Despite this, the total speedup remains 1.37 – 3.45$\times$ for the mPLUG-Owl3 models and 1.04 – 2.22$\times$ for the OpenFlamingo models.
@@ 4.3 Comparison with DeepSpeed-Ulysses
For Deepspeed-Ulysses, each attention operation involves two all-to-all communications: one before the computation to gather input query, key and value blocks, and another afterward to distribute attention output along the sequence dimension. The first all-to-all is expensive as it involves communicating the large key-value blocks. To see this, we compare Deepspeed-Ulysses with LV-XAttn on mPLUG-Owl3-2b using the cluster with A100 80GB GPUs. As shown in Table 5, LV-XAttn achieves 1.21 – 1.55$\times$ speedup compared to Deepspeed-Ulysses.
In addition, without activation recomputation, the larger memory footprint of Deepspeed-Ulysses limits its ability to process large visual inputs. When using the OpenFlamingo-3b model on the cluster with A30 24GB GPUs, Table 5 shows that LV-XAttn is able to process up to 4×4$\times$4 × longer text and visual inputs compared to Deepspeed-Ulysses.
@@ 4.4 Ablation Study
Activation Recomputation Figure 6(a) and 6(b) show the iteration time for running mPLUG-Owl-7b and OpenFlamingo-3b on a single node with 3 A100 40GB GPUs, with and without employing activation recomputation for cross-attention layers. By omitting the saving of large key-value blocks, the reduced memory consumption enables the processing of a larger number of frames, increasing by 1.6$\times$ and 1.5$\times$ for mPLUG-Owl-7b and OpenFlamingo-3b, respectively, with a negligible overhead of less than 8%.
@@ 5 Related Work
Multimodal Large Langauge Models There are two main classes of MLLM designs. The first design concatenates tokenized visual inputs with text tokens and feeds them into the LLM. Models such as LLaVA (Liu et al., 2023), InternVL (Chen et al., 2024), NVILA (Liu et al., 2024b), and MiniGPT-4 (Zhu et al., 2024) follow this paradigm. Although this approach naturally extends text-only LLMs, the large number of tokens significantly slows down both training and inference (Ye et al., 2024; Grattafiori et al., 2024). The second design relies on cross-attention mechanisms, where cross-attention layers are inserted between LLM layers to incorporate visual features into its intermediate representations. Models such as Flamingo (Alayrac et al., 2022), IDEFICS (Laurençon et al., 2024), Otter (Li et al., 2023a), mPLUG-Owl3 (Ye et al., 2024) and Llama 3-V (Grattafiori et al., 2024) adopt this strategy. While this method avoids processing a large number of visual tokens through the LLM backbone, cross-attention layers remain computationally expensive with current sequence-parallel approaches (Grattafiori et al., 2024). In this work, we introduce LV-XAttn to address this bottleneck.
@@
Parallelism Distributed attention can be classified into head-parallelism approaches, such as Megatron-LM (Korthikanti et al., 2023a) and Deepspeed-Ulysses (Jacobs et al., 2024), and sequence-parallelism approaches like Ring Attention, its variants DistFlashAttn (Li et al., 2024), and Striped Attention (Brandon et al., 2023). As discussed in Section 2, these methods face significant communication. Our work proposes a solution with minimal communication overhead for cross-attention in MLLMs. General parallelism approaches, such as data parallelism (Dean et al., 2012), pipeline parallelism (Narayanan et al., 2019), tensor parallelism (Shoeybi et al., 2019), and FSDP (Rajbhandari et al., 2020), can be combined with our approach.
@@ 6 Conclusion
We introduced LV-XAttn, a distributed, exact cross-attention mechanism for MLLMs with minimal communication overhead. By storing large key-value blocks locally on each worker and transmitting only smaller query blocks, LV-XAttn significantly reduces communication volume, which can be fully hidden by computation. Additionally, the activation recomputation technique reduces memory usage, enabling the processing of longer visual inputs with minimal overhead. Our evaluation demonstrates that LV-XAttn speeds up MLLM iteration by up to 5.58× and enables the processing of visual inputs up to 1.6× longer.
@@ Acknowledgements
The authors were supported by NSF award CNS 2237306. We also thank Zhao Zhang from Rutgers University for providing us access to computing resources. The authors acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award DDR-ERCAP0029980. This research also used computational resources from the NSF Cloudlab (Duplyakin et al., 2019) facility.
@@ Appendix A Comparison of LV-XAttn and Ring Attention for General Use Case
Figure 7 plots the theoretical speedup of LV-XAttn over Ring Attention for general $S_{Q}$ and $S_{KV}$. When $S_{KV}$ is large enough (above the horizontal bright red line), the transmission of $O_{i},Q_{i},m_{i}$ and $l_{i}$ in LV-XAttn is hidden by computation, making LV-XAttn compute-bound. On the other hand, when $S_{Q}$ is not too large (to the left of the vertical dark red line), the transmission of $K_{i}$ and $V_{i}$ are too large to be hidden by computation, making Ring Attention compute-bound. Their intersection (the top-left quadrant) represents the typical MLLM use case with large visual inputs and small text prompt.
For smaller visual inputs, the reduced $S_{KV}$ causes LV-XAttn to also become communication-bound (the bottom-left quadrant). When both LV-XAttn and Ring Attention are communication-bound, their relative speed depends on communication volume: LV-XAttn sends $3\cdot\frac{S_{Q}d}{n}+\frac{S_{Q}}{n}$, while Ring Attention sends $2\cdot\frac{S_{KV}}{n}$. Roughly, when $S_{KV}>\frac{3}{2}S_{Q}$, LV-XAttn still remains faster than Ring Attention. For MLLMs, each image is encoded into a large number of visual tokens – e.g., 729 for mPLUG-Owl3 and 64 for OpenFlamingo – so this condition is typically satisfied, making LV-XAttn faster for MLLMs in general.

### v3

@@ 4.1 Experimental Setup
Implementation LV-XAttn is implemented using PyTorch and Triton (Tillet et al., 2019). It uses torch.distributed for distributed communication, while the modified FlashAttention kernels to support rescaling operations are implemented with Triton. We validated the correctness of LV-XAttn by confirming that its output matches that of PyTorch’s scaled dot-product attention implementation.
@@
Model Setup We evaluate our methods and baselines on 6 models shown in Table 2. Using their publicly available model checkpoints, we replace the cross-attention operations with LV-XAttn and other distributed attention baselines. Following their default configuration, each frame is encoded into 6404 visual tokens for Llama 3-V, 729 visual tokens for the mPLUG-Owl3 models, and 64 visual tokens for the xOpenFlamingo models. This implies that given the same amount of memory capacity, OpenFlamingo models can accommodate more frames than Llama 3-V and mPLUG-Owl3 models.
@@
Baselines For our method, we use LV-XAttn for the cross-attention layers and Ring Attention for the LM blocks. Our primary baseline is the setup where Ring Attention is used for both the cross-attention layers and LM blocks. We apply our activation recomputation technique to both of these settings for enabling longer context length. We also compare against Deepspeed-Ulysses (Jacobs et al., 2024), which employs sequence parallelism for non-attention layers and head parallelism for attention layers. All methods use FlashAttention.
Our evaluation focuses on comparing the runtime of distributed cross-attention mechanisms. We follow the same benchmarking methodology used in previous work (Dao, 2024; Li et al., 2024), measuring runtime on randomly generated inputs of specific sizes to compare our method with the baselines across different scales. The reported runtime is averaged over 5 trials, following 2 warmup runs.
@@ 4.2 Comparison with Ring Attention
Table 3: Per iteration wall-clock time (in seconds) on 16 A100 80GB GPUs with Ring Attention and LV-XAttn. “CA” represents the time spent on cross-attention operations. As $S_{Q}$ doubles, the cross-attention speedup nearly halves because the runtime for Ring Attention, which is communication-bound, remains constant, while the runtime for LV-XAttn, which is computation-bound, doubles. On the other hand, as $S_{KV}$ doubles, both communication and computation also double, so the speedup remains roughly the same.
@@
Llama 3-V-11b 2K 384 2K 2401K 301.81 333.29 14.22 31.38 21.22$\times$ 10.62$\times$
2K 192 2K 1200K 144.5 167.2 8.36 22.27 17.28$\times$ 7.51$\times$
1K 192 1K 1200K 151.83 176.16 5.5 19.93 27.59$\times$ 8.84$\times$
@@
length count CA Total CA Total CA Total
Llama 3-V-11b 512 64 512 400K 109.07 154.18 3.35 49.24 32.56$\times$ 3.13$\times$
512 32 512 200K 55.8 101.13 2.49 45.9 22.43$\times$ 2.2$\times$
256 32 256 200K 57.5 100.75 1.7 44.93 33.77$\times$ 2.24$\times$
@@
Table 3 shows the per iteration time of six models using LV-XAttn and Ring Attention on 16 A100 80GB GPUs. For Llama 3-V, LV-XAttn speeds up the cross-attention operation by 17.28 – 27.59$\times$. Since the cross-attention operation accounts for the majority of the total iteration time when using Ring Attention, this reduction results in a significant total iteration speedup of 7.51 – 10.62$\times$. For the mPLUG-Owl3 and OpenFlamingo models, which process a larger number of frames (as each frame is encoded into fewer visual tokens compared to Llama 3-V) and thus have longer text lengths (due to the inclusion of a special token <image> per frame) and larger $S_{Q}$, the speedup of cross-attention remains significant, though less pronounced: 7.04 – 15.32$\times$ for the mPLUG-Owl3 models and 1.47 – 2.14$\times$ for the OpenFlamingo models. Notably, OpenFlamingo-3b, with denser cross-attention layers, spends a larger portion of its time in cross-attention compared to OpenFlamingo-9b when using Ring Attention. Consequently, the speedup in cross-attention translates to a more substantial end-to-end speedup for OpenFlamingo-3b.
Table 4 shows the same experiment on 8 A30 24GB GPUs. In this setup, we have smaller text lengths and fewer frames due to the smaller memory capacity. The speedup for cross-attention operation is greater than that on 16 A100 GPUs: 22.43 – 33.77$\times$ for Llama 3-V, 20.6 – 45.85$\times$ for the mPLUG-Owl3 models, and 3.97 – 7.2$\times$ for the OpenFlamingo models. This is due to smaller query block sizes $\frac{S_{Q}}{n}$ (shorter computations favors computation-bound LV-XAttn) and slower interconnect bandwidth (longer communication hurts communication-bound Ring Attention), as shown in Table 1. However, the larger cross-attention speedups do not translate into a larger total speedup, as the portion of time spent on cross-attention layers decreases due to slower self-attention layers in LM blocks (caused by lower GPU FLOPS and slower interconnect). Despite this, the total speedup remains 2.2 – 3.13$\times$ for Llama 3-V, 1.65 – 3.45$\times$ for the mPLUG-Owl3 models, and 1.04 – 2.22$\times$ for the OpenFlamingo models.
@@ 4.3 Comparison with DeepSpeed-Ulysses
For Deepspeed-Ulysses, each attention operation involves two all-to-all communications: one before the computation to gather input query, key and value blocks, and another afterward to distribute attention output along the sequence dimension. The first all-to-all is expensive as it involves communicating the large key-value blocks. To see this, we compare Deepspeed-Ulysses with LV-XAttn on mPLUG-Owl3-2b using the cluster with A100 80GB GPUs. As shown in Table 5, LV-XAttn achieves 1.34 – 1.55$\times$ speedup compared to Deepspeed-Ulysses.
In addition, without activation recomputation, the larger memory footprint of Deepspeed-Ulysses limits its ability to process large visual inputs. When running OpenFlamingo-3b on the cluster with A30 24GB GPUs, Table 6 shows that LV-XAttn is able to process more than 4×4$\times$4 × longer text and visual inputs compared to Deepspeed-Ulysses.
@@ 4.4 Ablation Study
Activation Recomputation Figures 6(a) and 6(b) show the iteration time for running mPLUG-Owl-7b and OpenFlamingo-3b on a single node with 3 A100 40GB GPUs, with and without employing activation recomputation for cross-attention layers. By omitting the saving of large key-value blocks, the reduced memory consumption enables the processing of a larger number of frames, increasing by 1.5$\times$ and 1.6$\times$ for mPLUG-Owl-7b and OpenFlamingo-3b, respectively, with a runtime overhead of less than 8%.
5 Discussion
In this section, we discuss the applicability of LV-XAttn to MLLMs with different architectures.
There are two main classes of MLLM architectures: concatenation-based and cross-attention-based. The first design concatenates tokenized visual inputs with text tokens and feeds them into the LLM backbone. Models such as LLaVA (Liu et al., 2023), InternVL (Chen et al., 2024), NVILA (Liu et al., 2024b), and MiniGPT-4 (Zhu et al., 2024) follow this paradigm. Concatenation-based models have shown strong unified multimodal reasoning capabilities. However, the large number of tokens fed to the LLM significantly slows down both training and inference, limiting their scalability and efficiency (Ye et al., 2024; Grattafiori et al., 2024).
The second design relies on cross-attention, where cross-attention layers are inserted between LLM layers to incorporate visual features into its intermediate representations. Models such as Flamingo (Alayrac et al., 2022), IDEFICS (Laurençon et al., 2024), Otter (Li et al., 2023a), mPLUG-Owl3 (Ye et al., 2024), Llama 3-V (Grattafiori et al., 2024), and NVLM-X (Dai et al., 2024) adopt this strategy. By bypassing the need to process a large number of visual tokens through the LLM backbone, cross-attention-based MLLMs offer superior computational efficiency when handling large visual inputs (Dai et al., 2024). For example, cross-attention-based MLLMs exhibit 10×10$\times$10 × better prefill latency than similar-sized concatenation-based MLLMs (Qiu et al., 2025).
However, as the size of visual inputs increase, even cross-attention-based MLLMs encounter efficiency bottlenecks due to communication overhead. To address this, we proposed LV-XAttn, which substantially reduces the communication volume to achieve significant speedup.
To achieve both high accuracy and computational efficiency, recent work have explored a hybrid architecture that uses both concatenation and cross-attention (Dai et al., 2024). We note that LV-XAttn can also be applied to such architecture to address communication bottlenecks.
@@ 6 Related Work
Parallelism Distributed attention can be classified into head-parallelism approaches, such as Megatron-LM (Korthikanti et al., 2023a) and Deepspeed-Ulysses (Jacobs et al., 2024), and sequence-parallelism approaches like Ring Attention, its variants DistFlashAttn (Li et al., 2024), and Striped Attention (Brandon et al., 2023). As discussed in Section 2, these methods face significant communication overhead. Our work proposes a solution with minimal communication overhead for cross-attention in MLLMs. General parallelism approaches, such as data parallelism (Dean et al., 2012), pipeline parallelism (Narayanan et al., 2019), tensor parallelism (Shoeybi et al., 2019), and fully sharded data parallelism (Rajbhandari et al., 2020), can be combined with our approach.
@@ 7 Conclusion
We introduced LV-XAttn, a distributed, exact cross-attention mechanism for MLLMs with minimal communication overhead. By storing large key-value blocks locally on each worker and transmitting only smaller query blocks, LV-XAttn significantly reduces communication volume, which can be fully hidden by computation. Additionally, the activation recomputation technique reduces memory usage, enabling the processing of longer visual inputs with minimal overhead. Our evaluation demonstrates that LV-XAttn speeds up MLLM iteration by up to 10.62$\times$ and enables the processing of visual inputs up to 1.6$\times$ longer.
@@ Acknowledgements
The authors were supported by NSF award CNS 2237306. We thank Zhao Zhang from Rutgers University for providing us access to computing resources. The authors acknowledge the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award DDR-ERCAP0029980. This research also used computational resources from the NSF Cloudlab (Duplyakin et al., 2019) facility. We also thank Wenxuan Tan for his insightful suggestions on improving the paper.
Impact Statement
This paper aims to improve the efficiency of MLLMs by alleviating communication bottlenecks when processing long visual inputs. By reducing end-to-end runtime, it helps lower the computational cost of training and deploying MLLMs, which in turn can reduce energy consumption and environmental impact. The approach does not alter model outputs, and thus does not introduce additional risks beyond those already associated with MLLMs.
@@ Appendix A Comparison of LV-XAttn and Ring Attention for General Use Case
Figure 7 plots the theoretical speedup of LV-XAttn over Ring Attention for general $S_{Q}$ and $S_{KV}$. When $S_{KV}$ is large enough (above the horizontal bright red line), the transmission of $Q_{i}$, $O_{i}$ and $L_{i}$ in LV-XAttn is hidden by computation, making LV-XAttn compute-bound. On the other hand, when $S_{Q}$ is small (to the left of the vertical dark red line), the transmission of $K_{i}$ and $V_{i}$ are too large to be hidden by computation, making Ring Attention compute-bound. Their intersection (the top-left quadrant) represents the typical MLLM use case with large visual inputs and small text prompt.
For smaller visual inputs, the reduced $S_{KV}$ causes LV-XAttn to also become communication-bound (the bottom-left quadrant). When both LV-XAttn and Ring Attention are communication-bound, their relative speed depends on communication volume: LV-XAttn sends $2\cdot\frac{S_{Q}d}{n}+\frac{S_{Q}}{n}$, while Ring Attention sends $2\cdot\frac{S_{KV}d}{n}$. Roughly, when $S_{KV}>S_{Q}$, LV-XAttn still remains faster than Ring Attention. For MLLMs, each image is encoded into a large number of visual tokens – e.g., 6404 for Llama 3-V, 729 for mPLUG-Owl3 and 64 for OpenFlamingo – so this condition is typically satisfied, making LV-XAttn faster for MLLMs in general.